# Protecting information in a parametrically driven hybrid quantum system

### v3

@@ I Introduction
A key aspect of working with an ensemble of two-level atoms or spin-$1/2$ particles is that the transition frequencies for each particle may not be identical, a feature known as inhomogeneous broadening. Such broadening has the detrimental effect of decohering any information stored or transmitted from the ensemble to a coupled cavity, thus limiting the performance of any information processing protocol [14]. Over the years, several approaches have been proposed and implemented to overcome the effect of inhomogeneous broadening. A very natural recourse is the effect of “cavity protection”, whereby the coupling of the spin-ensemble with the cavity is significantly increased to suppress the decoherence arising due to the broadening [15, 16, 17]. Spin-ensemble based hybrid systems already offer the advantage of operating in the strong-coupling regime, due to the enhanced collective coupling strength of a large number of spins [9, 8, 10]. However, increasing this coupling further to suppress decoherence may require operational regimes, such as high-Q cavities or increased spin density, which may either be difficult to implement or introduce unwanted spin interactions. Other approaches rely on refocusing mechanisms using spin-echo [18] or sophisticated engineering of the spectral distribution of the ensemble based on either optimal selection [19], dynamical decoupling [20] or hole-burning [21, 22]. While each method has its advantages, they often rely on complex control protocols or challenging experimental schemes to manipulate the spins, which may impede the performance of the hybrid system.
In this work, we propose a radically different method to overcome decoherence due to inhomogeneous broadening – one that does not involve altering the intrinsic spectral distribution of the spins but focuses on how information is encoded in the hybrid quantum system. In this approach, the quantum cavity is subjected to a parametric, two-photon drive, which exponentially enhances the coupling [23, 24] between the states of the cavity in the squeezed frame and the inhomogeneously broadened spin-ensemble. Therefore, any information now encoded in this new cavity frame experiences an implicit cavity protection effect. A clear advantage here is that by tuning the parametric driving strength, the effective coupling can be increased without resorting to the design of expensive cavities or the use of a high spin-density ensemble. Such enhancement in coupling has been studied in other contexts, including ultrastrong coupling between a few qubits and the cavity [25, 26], creation of entanglement [27, 28], superradiant phase transition [29, 30] and squeezed lasing [31].
@@
The paper is arranged in the following way. In Sec. II, we look at the model that describes the dynamics of the hybrid system consisting of an ensemble of spins inside a cavity under parametric driving. In Sec. III, we derive the semiclassical equations for the photon and spin excitation operators and study the transition from weak to strong coupling regime. For full quantum solutions, we present the variational renormalization group method for driven-dissipative systems in Sec. IV, and show the protection effect under increasing driving strength. Finally, we end with a discussion of the main results in Sec. V.
II The effective Hamiltonian
The dynamics of a hybrid quantum system, consisting of an ensemble of $N$ two-level atoms or spin-$1/2$ particles interacting with a quantum cavity, as shown in Fig. (1), is governed by the Tavis-Cummings model [36]. For a system subjected to a parametric, two-photon drive with a frequency $\omega_{d}$, the corresponding Hamiltonian ($\hbar=1$) under the rotating wave approximation (RWA) is given by,
$\mathcal{H}_{0}$ $=$ $\frac{1}{2}\sum_{k=1}^{N}\Delta_{k}\sigma_{k}^{z}+\Delta_{c}\hat{a}^{\dagger}\hat{a}+\sum_{k=1}^{N}g_{k}\left(\sigma_{k}^{-}\hat{a}^{\dagger}+\sigma_{k}^{+}\hat{a}\right)$ (1)
-- $\frac{\eta}{2}(\hat{a}^{2}+\hat{a}^{{\dagger}2}),$
where the system is in a frame rotating with half the driving frequency, $\omega_{d}/2$. The cavity and the spin transition frequencies are given by $\omega_{c}$ and $\omega_{k}$ for $k=1,\dots,N$. The corresponding frequency detunings with respect to $\omega_{d}/2$ are $\Delta_{c,k}=\omega_{c,k}-\omega_{d}/2$. As usual, $\hat{a}$ is the photon annihilation operator, $\{\sigma_{k}^{z},\sigma_{k}^{\pm}\}$ are the spin operators given by the Pauli matrices, $g_{k}$ is the coupling between the $k^{\textrm{th}}$ spin and the cavity field, and $\eta$ is the strength of the parametric drive.
@@ II The effective Hamiltonian
For this study, the fundamental quantity of interest is the inhomogeneity of the spin ensemble, given by the distribution of the spin transition frequencies $\omega_{k}$ and the spin-photon coupling $g_{k}$. The frequency distribution is taken to be Gaussian (cf. [38, 37]) with a standard deviation or width given by $\delta$. For simplicity, the coupling is taken to be identical i.e., $g_{k}=g,\forall~{}k$, which results in an effective coupling strength of $\Omega=\sqrt{\sum_{k=1}^{N}g_{k}^{2}}=\sqrt{N}g$. This demonstrates an enhancement of the coupling by a factor of $\sqrt{N}$ compared to that of a single spin. Hence, ensembles with a high number of atoms or spins readily exhibit phenomena related to strong light-matter coupling even when individual particles couple only weakly to the cavity [9, 10].
@@
Importantly, the effective spin-photon coupling can be further enhanced by parametrically driving the system. Using the unitary operator $\mathcal{U}=\exp\left[{r(a^{2}-a^{{\dagger}2})/2}\right]$, where $r=1/2\tanh^{-1}(\eta/\Delta_{c})$ is the squeezing parameter, the cavity states can be transformed to a squeezed frame. The transformed spin-ensemble-cavity Hamiltonian is then given by $\mathcal{H}_{sq}=\mathcal{U}\mathcal{H}_{0}\mathcal{U}^{\dagger}$, such that
@@
$\mathcal{H}_{sq}$ $=$ $\tilde{\Delta}_{c}\hat{a}_{s}^{\dagger}\hat{a}_{s}+\frac{1}{2}\sum_{k=1}^{N}\left\{\Delta_{k}\sigma_{k}^{z}+g_{k}e^{r}\left(\hat{a}_{s}+\hat{a}_{s}^{\dagger}\right)\left(\sigma_{k}^{-}+\sigma_{k}^{+}\right)\right.$ (2)
-- $\left.g_{k}e^{-r}\left(\hat{a}_{s}-\hat{a}_{s}^{\dagger}\right)\left(\sigma_{k}^{-}-\sigma_{k}^{+}\right)\right\},$
@@
where $\tilde{\Delta}_{c}=\Delta_{c}/\cosh{(2r)}$ and $\hat{a}_{s}=\cosh{(r)}\hat{a}+\sinh{(r)}\hat{a}^{\dagger}$ is the cavity operator in the squeezed frame. For large $r$ (i.e., $e^{-r}\rightarrow 0$), the Hamiltonian $\mathcal{H}_{sq}$ is nothing but the inhomogeneous quantum Dicke model or the $N$-spin Rabi model in the squeezed frame, with an exponentially enhanced coupling $\tilde{g}_{k}=g_{k}e^{r}/2$. Moreover, $\mathcal{H}_{sq}$ reduces to the Tavis-Cummings model in certain regimes where $|\tilde{\Delta}_{c}-\Delta_{k}|\ll\tilde{\Delta}_{c}+\Delta_{k}$ and $\tilde{g}_{k}\ll\tilde{\Delta}_{c}$, which ensures that RWA is valid for the transformed Hamiltonian and the system is not in the ultra-strong coupling regime [39, 40]. Therefore, the effective spin-cavity coupling in the transformed frame can be readily controlled simply by changing the parametric driving strength $\eta$.
Recently, such exponential enhancement in interaction strength was demonstrated in several experiments involving ion traps [41, 42, 43] and superconducting circuits [44]. In trapped-ion experiment [43], authors demonstrate amplification of interactions involving quantum harmonic oscillators, reaching squeezing as high as $r=1.38$ with typical parameter values of $\omega_{c}/2\pi\sim 7$ MHz and $g/2\pi\sim 1.5$ KHz. For superconducting circuit experiment [44], a resonator is capacitively coupled to a transmon qubit and authors show two-fold increase in dispersive coupling corresponding to $r\approx 0.81$ for $\Delta_{c}/2\pi=20$ MHz and $\eta/2\pi=17$ MHz. Another experiment demonstrates detection of $15$ dB squeezed vacuum states of light using degenerate parametric down conversion process. This corresponds to $r=1.73$ and is the highest squeezing observed till now [45]. In all of these experiments, various decoherence processes (including squeezing induced decoherences) adversely affects the overall dynamics of the system and limits the maximum squeezing that can be achieved.
III Dynamics in the semiclassical regime
@@
$\frac{d\rho}{dt}$ $=$ $-i[\mathcal{H},\rho]+\kappa\mathcal{L}_{\hat{a}}[\rho]+\sum_{k}^{n}\{\gamma_{h}\mathcal{L}_{\sigma^{-}_{k}}[\rho]+\gamma_{p}\mathcal{L}_{\sigma^{z}_{k}}[\rho]\},~{}~{}~{}$ (3)
where $\rho$ is the density matrix of the system and the Lindblad operators are $\mathcal{L}_{\hat{x}}[\rho]=\hat{x}\rho\hat{x}^{\dagger}-\frac{1}{2}\{\hat{x}^{\dagger}\hat{x},\rho\}$. The photon loss rate is $\kappa$, the rate of radiative decay and dephasing of spins is $\gamma_{h}$ and $\gamma_{p}$. The equations of motion for the photonic and spin excitations, $\langle\hat{a}\rangle$, $\langle\sigma_{k}^{-}\rangle$ and $\langle\sigma_{k}^{z}\rangle$ are calculated using Eq. (3), for the Hamiltonian $\mathcal{H}=\mathcal{H}_{0}$. Using semi-classical (or mean-field) approximation in the limit of $N\rightarrow\infty$, all correlations between the emitters and the cavity can be ignored. This allows factorisation of all first-order correlations, i.e., $\langle\sigma_{k}^{-}\hat{a}\rangle\approx\langle\sigma_{k}^{-}\rangle\langle\hat{a}\rangle$, in the equations of motion of the hybrid system and we get,
$\frac{d\braket{\hat{a}}}{dt}$ $=$ $-(\kappa+i{\Delta}_{c})\braket{\hat{a}}-i\sum_{k}g_{k}\braket{\sigma_{k}^{-}}+i\eta\braket{\hat{a}}^{*},$
$\frac{d\braket{\sigma_{k}^{-}}}{dt}$ $=$ $-(\gamma_{h}+2\gamma_{p}+i\Delta_{k})\braket{\sigma^{-}_{k}}+ig_{k}\braket{\sigma^{z}_{k}}\braket{\hat{a}},$
$\frac{d\braket{\sigma_{k}^{z}}}{dt}$ $=$ $-2\gamma_{h}(1+\braket{\sigma^{z}_{k}})+2ig_{k}(\braket{\sigma_{k}^{-}}\braket{\hat{a}}^{*}-\braket{\sigma_{k}^{+}}\braket{\hat{a}}),$
where, $\braket{a^{\dagger}}=\braket{a}^{*}$ and $\braket{\sigma_{k}^{+}}=\braket{\sigma_{k}^{-}}^{*}$. As the approximation ignores correlations between the ensemble and the cavity, it is not suited for studying quantum properties like entanglement and coherent exchange of information. However, it still allows the study of temporal evolution of average photon number and can even be used for large but finite $N$, away from criticality [33, 34].
By solving the equations above, we study the dynamics of a hybrid system consisting of an inhomogeneous ensemble of $N=10^{4}$ spins. This ensemble is sufficiently large to use the semi-classical approximation in the stable region [34]. The key parameters used in the calculation are cavity detuning $\Delta_{c}=70$ GHz, and the width of the spectral distribution, $\delta=50-80$ MHz with mean $\bar{\Delta}_{k}=\tilde{\Delta}_{c}$. The collective coupling is taken to be $\Omega=40$ MHz and the parametric driving strength $\eta=\Delta_{c}\tanh(2r)$, with squeezing parameter $r=0.0-2.4$. Also, to drive the system in linear regime, we ensure that $\tilde{g}_{k}\ll\tilde{\Delta}_{c}$ Moreover, to better characterize the effect of inhomogeneous broadening in the semiclassical limit, the photon and spin losses are taken to be negligible as compared to the coupling and width of the distribution. These parameters typically corresponds to NV centres, but the dynamics is performed in the units of the cavity detuning and therefore as long as proper rescaling is ensured, it remains independent of absolute parameter values.
Figure 2: Semiclassical dynamics of an ensemble of spins inside a cavity. The figures show (a) the time evolution of the average photon number $n(t)=|\braket{\hat{a}}_{sq}|^{2}$, for an ensemble with Gaussian frequency distribution of width $\delta=60$ MHz and mean $\bar{\Delta}_{k}=\tilde{\Delta}_{c}$, and squeezing parameter $r=0.0$ (red dashed), $r=1.0$ (blue dot-dashed), and $r=2.0$ (black dotted), and (b) the variation in decay rate $\zeta$ with $r$, for width $\delta=60$ MHz (red circle), $\delta=70$ MHz (blue cross), and $\delta=80$ MHz (black square). The error bars show the standard deviation between the actual value of $n(t)$ and those corresponding to the fitted decay rate $\zeta$.
Figure 2 captures the dynamics of a hybrid system under parametric driving. The semiclassical equations are derived using the Hamiltonian $\mathcal{H}_{0}$, such that the average photon number in the squeezed frame is given by $n(t)=|\braket{\hat{a}_{s}}|^{2}=\cosh(2r)|\braket{\hat{a}}|^{2}-\sinh(2r)\Re(\hat{a}^{2})$, where $r$ is the squeezing parameter. Similar results are obtained if one instead studies the equations of motion in the transformed frame by using $\mathcal{H}_{sq}$. In Fig. 2(a), the temporal evolution of $n(t)$ is shown for different values of squeezing parameter $r$, and a fixed width of the spin frequency distribution, $\delta=60$ MHz. It is assumed that the hybrid system is initialized such that the cavity has photon number equal to unity i.e., $n(0)=1$.
The figure shows that for $r=0$, the average photon number quickly decoheres with time due to the inhomogeneity in the spin ensemble (other sources of dissipation have been ignored at this point). However, these excitations are sustained for higher values of $r$, which shows that boosting the parametric drive leads to a reduced decay rate and a longer lifetime for the photon. The rate at which the cavity excitation $n(t)$ is decohering can be captured by fitting the peaks of the Rabi oscillations using the relation, $n(t)=n(0)e^{-\zeta\omega_{0}t}$, where $\zeta$ is a dimensionless quantity, and $\omega_{0}=10$ MHz is a reference frequency. So, higher $\zeta$ implies faster decoherence of the photonic excitations in the hybrid quantum system. Figure 2(b) shows the variation of the rate of decay $\zeta$ as a function of squeezing parameter $r$, for spin ensembles of different width $\delta$. The plots clearly show that loss of excitation and decoherence in the system is more severe for ensembles with more inhomogeneity (larger width $\delta$). Importantly, parametric drives of increasing intensities, which give us higher values of $r$, are able to reduce the decay rate $\zeta$ successfully even for significantly broadened spin ensembles, thus giving an overall “cavity protection” effect. In fact, the protection is also observed when the inhomogeneity is larger than the effective coupling, i.e., $\delta>\Omega$, albeit in the absence of any other dissipation in the system.
In presence of cavity and qubit decoherences, following the calculations in Ref. [15] for $\kappa,\gamma_{h},\gamma_{p}\ll\tilde{\Omega}$, it can be shown that the average photon number in the cavity decays at the rate, $\Gamma_{n}(\tilde{\Omega})=\kappa+\gamma_{h}+2\gamma_{p}+\pi\rho(\tilde{\Omega})\tilde{\Omega}^{2}$, where for Gaussian distribution of emitters, $\rho(\tilde{\Omega})\sim 2^{-\tilde{\Omega}^{2}/\delta^{2}}/\delta$ governs the decoherence induced by inhomogeneous broadening and $\tilde{\Omega}=\Omega\cosh{(r)}$ is the effective coupling strength in the squeezed frame. For $\tilde{\Omega}\ll\delta$, due to Purcell effect, the decay rate increases on increasing $r$ [16]. However, in strong coupling regime when $\tilde{\Omega}\gg\delta$, there is a protective energy gap between the superradiant and the subradiant spin-wave modes, which does not allow inhomogeneity to induce decoherence in the system [14]. As a result, an exponential decay in the decay rate $\Gamma_{n}$ is observed and eventually the dynamics of the system becomes independent of inhomogeneous broadening. To quantify the suppression in errors due to inhomogeneity, we define the ratio,
$\xi_{r}\equiv\frac{\Gamma_{n}(\tilde{\Omega})}{\Gamma_{n}(\infty)}-1=\frac{\pi\rho(\tilde{\Omega})\tilde{\Omega}^{2}}{\kappa+\gamma_{h}+2\gamma_{p}},$ (5)
which becomes zero in the limit of $r\rightarrow\infty$.
IV Protecting information in the hybrid system
While semiclassical equations highlight the revival of macroscopic quantities such as the average photon number in the cavity, to study protection of information encoded in any superposed quantum state, the dynamics of the full quantum system is necessary. This is a computationally challenging task due to the large Hilbert space of the spin-ensemble-cavity system. To overcome this, we use a variational renormalization group method to study driven-dissipative dynamics [35], which is a tensor-network method similar to density matrix renormalization group and matrix-product states [46]. The method employs a time-adaptive approach to temporally evolve the density matrix $\rho$ of a spin-cavity system of about hundred spins, using the Lindblad ME formalism. To achieve this, all states $\rho$ and operators $\hat{O}$ are mapped to a higher dimensional space as superkets $|\rho\rangle$ and superoperators $\hat{\mathcal{O}}$ i.e., $\rho\rightarrow\textbf{vec}(\rho)=|\rho\rangle$ and $\hat{O}\rho\rightarrow(\hat{O}\otimes\mathbb{I})|\rho\rangle=\hat{\mathcal{O}}|\rho\rangle$. This allows the master equation in Eq. (3) to be mapped to a Schrödinger like equation, $d|\rho\rangle/dt=\tilde{\mathcal{L}}|\rho\rangle$, in the superoperator space, where
~$\tilde{\mathcal{L}}=-i(\mathcal{H}\otimes\textbf{I}-\textbf{I}\otimes\mathcal{H}^{T})+\kappa\tilde{\mathcal{L}}_{\hat{a}}+\sum_{k}\{\gamma_{h}\tilde{\mathcal{L}}_{\sigma_{k}^{-}}+\gamma_{p}\tilde{\mathcal{L}}_{\sigma_{k}^{z}}\},$ (6)
and ~$\tilde{\mathcal{L}}_{\hat{x}}=\hat{x}\otimes\hat{x}^{*}-\frac{1}{2}\hat{x}^{\dagger}\hat{x}\otimes\textbf{I}-\frac{1}{2}\textbf{I}\otimes\hat{x}^{T}\hat{x}$. At this point, the variational renormalization has two key parts. First, finding the renormalized representation for the superket $|\rho\rangle$ in a significantly reduced subspace, and second, evolving the initial state $|\rho\rangle$ in a time-adaptive manner. As with most tensor-network methods, the renormalized space is obtained by eliminating the null space or those with marginal singular values. The time evolution is then governed using the Lindblad ME in Eq. (6), using the transformed Hamiltonian $\mathcal{H}=\mathcal{H}_{sq}$, and evolved adaptively for small intervals $\Delta t$, such that |$\ket{\rho(t+\delta t)}=e^{\tilde{\mathcal{L}}\Delta t}\ket{\rho(t)}$. Now, ~$\tilde{\mathcal{L}}=\sum_{k}\tilde{\mathcal{L}}_{k}$, and the second-order Suzuki-Trotter decomposition helps expand $e^{\tilde{\mathcal{L}}2\Delta t}$ into a product of single spin-cavity evolution: $e^{\tilde{\mathcal{L}}_{N}\Delta t}e^{\tilde{\mathcal{L}}_{N-1}\Delta t}\dots e^{\tilde{\mathcal{L}}_{1}2\Delta t}\dots e^{\tilde{\mathcal{L}}_{N-1}\Delta t}e^{\tilde{\mathcal{L}}_{N}\Delta t}$ (see Refs. [35, 47]). The computation then follows an iterative sweeping protocol as used in other time-adaptive tensor-network methods [48, 49].
We note that squeezing the cavity also transforms the Lindblad operator $\mathcal{L}_{\hat{a}}$, which is given by [25],
$\mathcal{L}_{\hat{a}}[\rho]\rightarrow(n_{\text{th}}+1)\mathcal{L}_{\hat{a}_{s}}[\rho]+n_{\text{th}}\mathcal{L}_{\hat{a}_{s}^{\dagger}}[\rho]\\ -m_{\text{2-ph}}\mathcal{L}^{\prime}_{\hat{a}_{s}}[\rho]-m_{\text{2-ph}}^{*}\mathcal{L}^{\prime}_{\hat{a}_{s}^{\dagger}},$ (7)
where $\mathcal{L}^{\prime}_{\hat{x}}[\rho]=\hat{x}\rho\hat{x}-\frac{1}{2}\{\hat{x}^{2},\rho\}$ and $n_{\text{th}}$, $m_{\text{2-ph}}$ are functions of squeezing parameter $r$, which correspond to thermal noise and two-photon correlations introduced by squeezing, respectively. This effectively induces additional errors in the system, but these can be eliminated by coupling the cavity mode to a squeezed-vacuum reservoir with the same squeezing $r$, which is phase-shifted by $\pi$ relative to the two-photon drive.
Figure 3: Dynamics of quantum information encoded in a hybrid system consisting of $N=100$ spins inside a cavity. The width of the frequency distribution is $\delta=30$ MHz. The plots in (a) show the evolution of the fidelity $\mathcal{F}$ between the effective cavity state at time $t$ and the initial state at $t=0$, for different $r$. The plot is the best-fit curve that envelopes the fidelity. The dashed curve corresponds to the fidelity averaged over 100 Haar random initial states for $r=2.0$. The other subfigures show the Wigner function $W(p,q)$ of the initial cavity state in (b), and the state at $t$ corresponding to fidelity (black triangles in (a)) for parameters (c) $r=0.0$ and (d) $r=2.0$.
For the full quantum dynamics of the hybrid system using variational renormalization group, we consider an ensemble of $N=100$ spins inside the cavity. For consistency with semiclassical results, we use the same set of parameters, including the collective coupling strength $\Omega=40$ MHz. However, we now consider other dissipative terms, such as the photon loss rate $\kappa=7$ MHz, and radiative loss and dephasing for spins equal to $\gamma_{h}=\kappa/8$ and $\gamma_{p}=\kappa/16$, respectively.

### v4

@@ 1 Introduction
A key aspect of working with an ensemble of two-level atoms or spin-$1/2$ particles is that the transition frequencies for each particle may not be identical, a feature known as inhomogeneous broadening. Such broadening has the detrimental effect of decohering any information stored or transmitted from the ensemble to a coupled cavity, thus limiting the performance of any information processing protocol [14]. Over the years, several approaches have been proposed and implemented to overcome the effect of inhomogeneous broadening. A natural recourse is the “cavity protection effect”, whereby the coupling between the spin-ensemble and the cavity is significantly increased to suppress the decoherence arising due to broadening [15, 16, 17, 18, 19]. Spin-ensemble based hybrid systems already offer the advantage of operating in the strong-coupling regime, due to the enhanced collective coupling strength of a large number of spins [8, 9, 10]. However, increasing this coupling further to suppress decoherence may require operational regimes, such as high-Q cavities or increased spin density, which may either be difficult to implement or introduce unwanted spin interactions. Other approaches rely on refocusing mechanisms using spin-echo [20] or sophisticated engineering of the spectral distribution of the ensemble based on either optimal selection [21], dynamical decoupling [22] or hole-burning [23, 24]. While each method has its advantages, they often rely on complex control protocols or challenging experimental schemes to manipulate the spins, which may impede the performance of the hybrid system.
In this work, we propose a radically different method to overcome decoherence due to inhomogeneous broadening – one that does not involve altering the intrinsic spectral distribution of the spins but focuses on how information is encoded in the hybrid quantum system. In this approach, the quantum cavity is subjected to a parametric, two-photon drive, which exponentially enhances the coupling [25, 26] between the states of the cavity in the squeezed frame and the inhomogeneously broadened spin-ensemble. Therefore, any information now encoded in this new cavity frame experiences an implicit cavity protection effect. A clear advantage here is that by tuning the parametric driving strength, the effective coupling can be increased without resorting to the design of expensive cavities or the use of a high spin-density ensemble. Such enhancement in coupling has been studied in other contexts, including ultrastrong coupling between a few qubits and the cavity [27, 28], creation of entanglement [29, 30], superradiant phase transition [31, 32] and squeezed lasing [33]. Further, recent experiments have demonstrated such enhancement in different platforms, including trapped ions [34, 35, 36] and superconducting circuits [37].
@@
The paper is arranged in the following way. In Sec. 2, we look at the model that describes the dynamics of the hybrid system consisting of an ensemble of spins inside a cavity under parametric driving. In Sec. 3, we derive the semiclassical equations for the photon and spin excitation operators and study the transition from weak to strong coupling regime. For full quantum solutions, we present the variational renormalization group method for driven-dissipative systems in Sec. 4, and show the protection effect under increasing driving strength. Finally, we end with a discussion of the main results in Sec. 5.
2 The effective Hamiltonian
The dynamics of a hybrid quantum system, consisting of an ensemble of $N$ two-level atoms or spin-$1/2$ particles interacting with a quantum cavity, as shown in Fig. (1), is governed by the Tavis-Cummings model [42]. For a system subjected to a parametric, two-photon drive with a frequency $\omega_{d}$, the corresponding Hamiltonian ($\hbar=1$) under the rotating wave approximation is given by,
$\mathcal{H}_{0}=\frac{1}{2}\sum_{k=1}^{N}\Delta_{k}\hat{\sigma}_{k}^{z}+\Delta_{c}\hat{a}^{\dagger}\hat{a}\\ +\sum_{k=1}^{N}g_{k}\left(\hat{\sigma}_{k}^{-}\hat{a}^{\dagger}+\hat{\sigma}_{k}^{+}\hat{a}\right)-\frac{\eta}{2}(\hat{a}^{2}+\hat{a}^{{\dagger}2}),$ (1)
where the system is in a frame rotating with half the driving frequency, $\omega_{d}/2$. The cavity and the spin transition frequencies are given by $\omega_{c}$ and $\omega_{k}$ for $k=1,\dots,N$. The corresponding frequency detunings with respect to $\omega_{d}/2$ are $\Delta_{c,k}=\omega_{c,k}-\omega_{d}/2$. As usual, $\hat{a}$ is the photon annihilation operator, $\{\hat{\sigma}_{k}^{z},\hat{\sigma}_{k}^{\pm}\}$ are the spin operators given by the Pauli matrices, $g_{k}$ is the coupling between the $k^{\textrm{th}}$ spin and the cavity field, and $\eta$ is the strength of the parametric drive.
@@ 2 The effective Hamiltonian
For this study, the fundamental quantity of interest is the inhomogeneity of the spin-ensemble, given by the distribution of the spin transition frequencies $\omega_{k}$ and the spin-photon coupling $g_{k}$. The frequency distribution is taken to be Gaussian [43, 44] (cf. [45]) with a standard deviation or width given by $\delta$. For simplicity, the coupling is taken to be identical i.e., $g_{k}=g,\forall\leavevmode\nobreak\ k$, which results in an effective coupling strength of $\Omega=\sqrt{\sum_{k=1}^{N}g_{k}^{2}}=\sqrt{N}g$. This demonstrates an enhancement of the coupling by a factor of $\sqrt{N}$ compared to that of a single spin. Hence, ensembles with a high number of atoms or spins readily exhibit phenomena related to strong light-matter coupling even when individual particles couple only weakly to the cavity [8, 10].
@@
Importantly, the effective spin-photon coupling can be further enhanced by parametrically driving the system. Using the unitary operator $\mathcal{U}=\exp\left[{r(\hat{a}^{2}-\hat{a}^{{\dagger}2})/2}\right]$, where $r=1/2\tanh^{-1}(\eta/\Delta_{c})$ is the squeezing parameter, the cavity states can be transformed to a squeezed frame. The transformed spin-ensemble-cavity Hamiltonian is then given by $\mathcal{H}_{sq}=\mathcal{U}\mathcal{H}_{0}\mathcal{U}^{\dagger}$, such that
@@
$\mathcal{H}_{sq}=\tilde{\Delta}_{c}\hat{a}_{s}^{\dagger}\hat{a}_{s}+\frac{1}{2}\sum_{k=1}^{N}\left\{\Delta_{k}\hat{\sigma}_{k}^{z}\right.\\ +g_{k}e^{r}\left(\hat{a}_{s}+\hat{a}_{s}^{\dagger}\right)\left(\hat{\sigma}_{k}^{-}+\hat{\sigma}_{k}^{+}\right)\\ -\left.g_{k}e^{-r}\left(\hat{a}_{s}-\hat{a}_{s}^{\dagger}\right)\left(\hat{\sigma}_{k}^{-}-\hat{\sigma}_{k}^{+}\right)\right\},$ (2)
@@
where $\tilde{\Delta}_{c}=\Delta_{c}/\cosh{(2r)}$ and $\hat{a}_{s}=\cosh{(r)}\hat{a}+\sinh{(r)}\hat{a}^{\dagger}$ is the cavity operator in the squeezed frame. For large $r$, such that $e^{-r}\to 0$, the Hamiltonian $\mathcal{H}_{sq}$ is nothing but the quantum Dicke model or the $N$-spin Rabi model with inhomogeneous spins and an exponentially enhanced coupling $\tilde{g}_{k}=g_{k}e^{r}/2$. The system is in resonance if the mean spin detuning $\bar{\Delta}_{k}=\sum_{k}\Delta_{k}/N\approx\tilde{\Delta}_{c}$. For $\tilde{g}_{k}\approx\tilde{\Delta}_{c}$, the system is in the ultra-strong coupling regime [46, 47]. However, for $\tilde{g}_{k}\ll\tilde{\Delta}_{c}$, the rotating wave approximation is valid and $\mathcal{H}_{sq}$ reduces to the Tavis-Cummings model with collective coupling $\tilde{\Omega}=\Omega\cosh{(r)}$. As such, the system can be studied in different operational regimes by manipulating the effective spin-cavity coupling, which can be readily controlled by changing the parametric driving amplitude $\eta$.
3 Dynamics in the semiclassical regime
@@
$\frac{d\rho}{dt}=-i[\mathcal{H},\rho]+\kappa\mathcal{L}_{\hat{a}}[\rho]\\ +\sum_{k=1}^{N}\{\gamma_{h}\mathcal{L}_{\hat{\sigma}^{-}_{k}}[\rho]+\gamma_{p}\mathcal{L}_{\hat{\sigma}^{z}_{k}}[\rho]\},$ (3)